Grade 12Physical Chemistry

alpha particle,electron and proton have equal kinetic energy.which particle has more ionizing power?

Anu
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
That one which woukd have greater velocity.
From equal K.E we get,
(2mp + 2me)v1^2 =mpv2^2 = mev3^2
here mass of electron is least and thereore it would have greater velocity and hence the elctron could have the more ionizing power.